OpenStudy (anonymous):

I need help in calculus. how would you go about getting the 1000th derivative of xe^-x

OpenStudy (anonymous):

\[f(x)=xe ^{-x}\]

OpenStudy (freckles):

I would find the first few derivatives and see if I could spot a pattern amongst them

OpenStudy (anonymous):

\[f'(x)=e ^{-x}-xe ^{-x}\]

OpenStudy (freckles):

\[f^{(1)}=e^{-x}(1-x) \\ f^{(2)}=e^{-x}(-1)+-e^{-x}(1-x)=-e^{-x}(1+1-x)=-e^{-x}(2-x)\] find the third one and put it in this form (factored form)

OpenStudy (freckles):

I think three will be all you need to see the pattern but you can do the 4th one if you are still not sure

OpenStudy (anonymous):

Did I make a mistake in the third derivative I got: \[f'''(x)=2e ^{x}-e ^{-x}+xe ^{-x}\]

OpenStudy (anonymous):

oh wait I made a mistake in my second derivative which caused my third derivative to incorrect

OpenStudy (freckles):

you don't have to multiple it out then find the derivative

OpenStudy (anonymous):

ok one second

OpenStudy (anonymous):

\[f'''(x)=e ^{-x}(3-x)\]

OpenStudy (freckles):

beauty

OpenStudy (anonymous):

awesome

OpenStudy (freckles):

\[f^{(1)}=e^{-x}(1-x) \\ f^{(2)}=-e^{-x}(2-x) \\ f^{(3)}=e^{-x}(3-x) \\ .... \\ f^{(1000)}=?\]

OpenStudy (anonymous):

so the 1000th derivative just be \[= e ^{-x}(1000-x)\]

OpenStudy (freckles):

almost!

OpenStudy (freckles):

you have one more thing to attach to that

OpenStudy (anonymous):

oh negative since its even

OpenStudy (freckles):

yep

OpenStudy (anonymous):

ah I see

OpenStudy (freckles):

like if you did go ahead and find the 4th derivative you would see something similar happen that happened in the 2nd derivative

OpenStudy (freckles):

\[f^{(4)}=(e^{-x})'(3-x)+(3-x)'(e^{-x}) \\ f^{(4)}=(-e^{-x})(3-x)+(0-1)(e^{-x}) \\ f^{(4)}=(-1)e^{-x}(3-x)+(-1)e^{-x} \\ f^{(4)}=-e^{-x}(3-x+1) \\ f^{(4)}=-e^{-x}(4-x)\]

OpenStudy (freckles):

it should be pretty easy to come up with the nth derivative

OpenStudy (anonymous):

right so theres a pattern, if the derivative is a even number then there is a negative infront of it.

OpenStudy (freckles):

\[f^{(n)}=(-1)^{n+1}e^{-x}(n-x)\]

OpenStudy (anonymous):

yup that makes sense, Thank you! :)

OpenStudy (freckles):

np